Arjhay and Equanne are travelling in the same direction, at the same speed of 100 kph on a level road. Arjhay is 150 meters ahead of Equanne. Arjhay decelerates at the rate of 3m/. From this instant up to the time Equanne overtakes Arjhay, how far Equanne traveled?


1
Expert's answer
2021-02-22T07:13:18-0500

First, calculate the distance it will take for Arjhay to decelerate from 100 kph to 0:


"d=\frac{v_1^2-v_2^2}{2a}=128.6\text{ m}."

So, Equanne traveled 150+128.6=278.6 m.